silemus.

7Binomial P(X≥1)P(X \ge 1)

X∼Bin(4,0.2)\displaystyle X \sim \mathrm{Bin}(4, 0.2). Find P(X≥1)\displaystyle P(X \ge 1). Give the answer to 4 decimal places.

At least one success is easier found as 1 minus the chance of zero successes. Here P(X=0) = 0.8⁴ = 0.4096, so P(X≥1) = 1 − 0.4096 = 0.5904.

How good is this answer?